02/08/10 – Payapp has arrived. TEI finished the as-builts drawings. We are shipping the original record markup dwgs, cd’s and final payapp to Bert for closeout.

02/04/10 – Look i can’t take it any longer. This project has to be moved to the Archive. It is just too old. I will update the final QCE payapp when it arrives. It is being moved today.

01/28/09 – TEI contacted QCE but Cavanaugh is out. We have not received their final payapp. As-Builts are almost done with Cad markups by TEI.

12/31/09 -QCE turned in closeout documents today for approval to TEI. TTC received and approved. As-Builts will be updated in 2 weeks by TEI. We need a final payapp from Queen City still for approval which is coming next week according to Mike C. After that this job will be closed out.

09/22/09 – We have contacted Queen City several times about closeout and still have not received the documents.

08/20/09 – QCE payapp has been approved but we have not received their closeout documents yet.

07/29/09 – QCE reported today that they know the O&M’s are due but they have just been too busy on trying to get new work in today’s market to finish them up. TTC told them that they could bill for retainage on this project and leave the closeout document line item at 0% until they submit the O&M’s. Cavanaugh said he might do this but they will work on the closeout docs.

06/23/09 – The final EDI item is the revised sequence which TEI wrote and sent to Data Aire and EDI for approval. The sequence has been approved but TTC is waiting in equipment deadband from DA. This is not critical to finishing the sequence. TTC may just update and send final to Bert. All other project contractors are complete. Get an update on closeout from Bert on EDI, QCE, CDFP paperwork.

05/22/09 – Data Aire Unit is running as designed on both chilled water and DX cooling. The system is changing over quickly and keep the space cool. TEI to write final sequence for the project.

04/15/09 – CR Architecture and Design awaiting final punchlist items to be completed.

04/03/09 – All work and punchlist items have been completed by EDI on this project. O&M’s have been logged in. Final payapp has been received. The payapp #2 was on hold but has now been release to Bert. The CDFP closeout paperwork was signed and sent to Bert yesterday. Queen City still owes us final payapp and O&M’s on this project. EDI still has to use chillers in the building before we know everything works.

02/18/09 – A controls meeting is scheduled for Friday with EDI and Data Aire.

02/07/09 – TTC requested a revised payment from EDI for this project.

02/02/09 – Terry & Mike will be down for project punchlist on Feb 4 for final walk through. The new insulation worked perfect through the cold snap last week. Bert reported no condensation on the unit, even on the the through metal parts.

01/22/09 – EDI is at the site today installing the insulation inside the unit today. All interior surfaces.

12/20/08 – Problems with the Data Aire Unit still exist. Condensation is a real problem. The unit is not insulated. DA said they can fix this problem with insulation. The controls and the chilled water pump are still not interlocked. A JCI N2 card can help the situation. Duft said that would be the easiest to use. TTC said it was going to be provided at no cost from DA on this project.

11/25/08 – TEI is working feverishly with EDI and QCE for final change orders. Hours were spent trying to resolve QCE change order #3 today and it still wasn’t exactly right. Another $317 needs to be written for QCE. ThermalTech must complete final punchlist on this project.

11/24/08 – Project is 95% complete. Final Change Orders need to be in by 11/26. TEI is working on them today.

10/16/08 – Bert reports that Stewart will be making all final computer moves this weekend and by Monday all essential Auditor equipment will be on the 5th floor. Demo of the 4th could start shortly for all contractors. QCE has made more modifications to the plug strip on the server racks and disconnected the 4th floor rack to facilitate its move to the 5th floor. The project is finally making progress towards completion.

09/18/08 – Trying to setup TCO for next week.

09/17/08 – Server must be running by next Tuesday. Fan test not complete. Temporary route condensate to 4th floor hub drain. EDI ductwork not installed. Power outage has delayed final fabrication of ductwork.

09/04/08 – Elevator quote came into today for $3k. Work to be done next Monday with Thyssen Krupp and EDI.

08/2508 – TTC sent out the plumbing Bulletin drawings to Helmes plumbing et al.

08/22/08 – TTC visited the site today for floor drain relocation. UPS is in and being wired. Disconnects are installed. Fire protection is in 5th floor room.

08/08/08 – Const Mtg notes from today include

08/01/08 – GM located new locations for the bus plugs disconnects for the project. One on the 5th and one on the 4th. QCE will have to move an existing disconnect to make the new ones fit.

07/30/08 – Data Aire Unit ship date is August 20. It will arrive 3 or 4 days after that. Quick ship prices were given but overly costly to the project for the minimal gain.

07/17/08 – Const Mtg. We discussed project schedule for this project. Contractors reviewed progress. Demo found a stairwell pressurization fan in the new computer room that must remain as is. EDI will seal ductwork. Computer room unit delivery is 4 to 5 weeks.

05/30/08 – Pre Const Mtg complete. Equip 5 weeks out. Contractors to provide schedule for start of the project.